.A transformer that its number of turns in the primary coils is 200 turns, and 400 tuns in the secondary coil, if the supplied voltage difference is 10V, calculate the secondary voltage difference?
1
Expert's answer
2020-03-18T10:27:31-0400

Applying the formula "\dfrac{N_P}{N_S}=\dfrac{V_P}{V_S}"

where Np is number of turns in the primary coil and Ns is number of turns in the secondary coil

and Vp is the voltage difference in the primary coil and Vs is the voltage difference in the secondry coil

"\dfrac{200}{400}=\dfrac{10}{V_S}\\V_S=20 volts"